Badger Valve and Fitting Company, located in southern Wisconsin, manufactures a variety of industrial valves and pipe fittings that are sold to customers in nearby states. Currently, the company is operating at about 70 percent capacity and is earning a satisfactory return on investment. Management has been approached by Glasgow Industries Ltd. of Scotland with an offer to buy 100,000 units of a pressure valve. Glasgow Industries manufactures a valve that is almost identical to Badger’s pressure valve; however, a fire in Glasgow Industries’ valve plant has shut down its manufacturing operations. Glasgow needs the 100,000 valves over the next four months to meet commitments to its regular customers. Glasgow is prepared to pay $30.20 each for the valves. Badger’s total product cost, based on current attainable standards, for the pressure valve is $31.50, calculated as follows:

Direct material $ 8.00
Direct labor 9.50
Manufacturing overhead 14.00
Total product cost $ 31.50

Manufacturing overhead is applied to production at the rate of $28 per standard direct-labor hour. This overhead rate is made up of the following components.

Variable manufacturing overhead $ 9.00
Fixed manufacturing overhead (traceable) 13.00
Fixed manufacturing overhead (allocated) 6.00
Applied manufacturing overhead rate $ 28.00

Additional costs incurred in connection with sales of the pressure valve include sales commissions of 5 percent and freight expense of $1.30 per unit. However, the company does not pay sales commissions on special orders that come directly to management. In determining selling prices, Badger adds a 30 percent markup to total product cost. This provides a $40.95 suggested selling price for the pressure valve. The Marketing Department, however, has set the current selling price at $39.45 in order to maintain market share. Production management believes that it can handle the Glasgow Industries order without disrupting its scheduled production. The order would, however, require additional fixed factory overhead of $15,000 per month in the form of supervision and clerical costs. If management accepts the order, 25,000 pressure valves will be manufactured and shipped to Glasgow Industries each month for the next four months. Glasgow’s management has agreed to pay the shipping charges for the valves.

1.Determine how many direct-labor hours would be required each month to fill the Glasgow Industries order

2.Prepare an analysis showing the impact of accepting the Glasgow Industries order. (Round "Per unit" answers to 2 decimal places.)

Per Unit Totals for 100,000 Units
Incremental revenue
Incremental costs:
Variable costs:
Direct material
Direct labor
Variable overhead
Total variable costs $0.00 $0
Fixed overhead:
Supervisory and clerical costs
Total incremental costs $0
Total incremental profit $0

3.Calculate the minimum unit price that Badger Valve and Fitting Company’s management could accept for the Glasgow Industries order without reducing net income. (Round your answer to 2 decimal places.)
Answer #1
1 Direct labor hours required for each month=Pressure valves to be produced per month*Labor hours required per valve
Labor hours required per valve=Applied manufacturing overhead rate/Manufacturing overhead in total product cost=28/14=0.50
Direct labor hours required for each month=25000*0.50=12500 hours
2 Per unit Totals for 100000 units
Incremental revenue (100000*30.2) 30.2 3020000
Incremental costs:
Variable costs:
Direct material (100000*8) 8 800000
Direct labor (100000*9.5) 9.5 950000
Variable overhead (100000*4.5) 4.5 450000
(9*0.5)
Total variable costs 22 2200000
Fixed overhead:
Supervisory and clerical costs (15000*4) 60000
Total incremental costs 2260000
Total incremental profit 760000
3 Minimum unit price=Variable costs per unit+Additional fixed cost per unit
Additional fixed costs per unit=Additional fixed cost/Pressure valves produced=60000/100000=$ 0.6 per valve
Minimum unit price=22+0.6=$ 22.6
I appreciate your ratings